consider a parallel plate capacitor with value of capacitance C =100(10^-6) F, the plate area,A =0.18m^2 and dielectric constant, εr=7.0
1
Expert's answer
2020-01-03T09:19:20-0500

A parallel plate capacitor with a dielectric between its plates has a capacitance given by "C=k \dfrac{A\epsilon_0}{d}" , where κ is the dielectric constant of the material,"A" is the area of the plate and "d" is the distance between the two plates and "\epsilon_0" is called the permittivity of free space whose value is "8.85*10^{-12}"

putting all the values in the above equation


"\implies100*10^{-6}=7*0.18*8.85*10^{-12}/d"

"\implies d=0.11*10^{-6} m"

"\implies" "d=0.11{\mu}m"